Merone Company allocates materials handling cost to the company's two products using the below data:

Modular Homes Prefab Barns
Total expected units produced 5,900 8,900
Total expected material moves 590 190
Expected direct labor-hours per unit 790 290

The total materials handling cost for the year is expected to be $325,890.

If the materials handling cost is allocated on the basis of material moves, the total materials handling cost allocated to the modular homes is closest to: (Round your intermediate calculations to 2 decimal places.)

Multiple Choice

  • $246,507.90

  • $238,382.50

  • $160,629.00

  • $213,514.00
Answer #1
Answer is $246,507.90
Statement showing Computations
Particulars Amount
total materials handling cost         325,890.00
Total expected material moves = 590 + 190                780.00
Cost per material moves = 325,890/780                417.81
total materials handling cost allocated to the modular homes is closest to = 417.81*590       246,507.90